Tohban Report 2014-06-03

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 28 May 2014
End Date: 03 Jun 2014
Tohban: Nicole Duncan (Vice Tohban: Ian Hannah)
Tohban email:
Next Tohban: Mitsuo (Vice Tohban: Hugh)
List all reports


Solar Activity

The sun was very quiet this past week. The big event was a small M-class flare on June 02. Note that although this flare was registered as M-class in GOES, hsi_tohban_flare_counter did not recognize an M-class event during this time period. The active region that produced this M and 2 C events, 12077, is near disk-center and will be on disk for the next week. The remaining 4 ARs on disk have had no flaring activity. By late in the next week, an AR should rotate on disk.

How many GOES flares occurred?

 Flares above B, C, M, X class were      11    8     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      8    8     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      5    3     0     0

There were RHESSI flares/GOES flares 47 / 19 over the time range 27-May-14 03-Jun-14

Memory Management

The SSR is emptying at the end of the pass set, with a max fill of 31%. It is emptying just at the end of the pass set.

We're still in Active/Vigorous decimation. Eventhough we have relatively little solar activity, the SSR is still filling with detector noise counts.

Spacecraft Status

The cold plate 1 and 2 are at 118.9 K and 117.4 K respectively. Spin rate 14.8.

Data Gaps



2014-06-02T12:15:00.000 -- 2014-06-02T12:20:00.000 300.00000

2014-05-29T04:50:00.000 -- 2014-05-29T05:00:00.000 600.00000

2014-05-29T06:25:00.000 -- 2014-05-29T06:30:00.000 300.00000

2014-05-29T06:45:00.000 -- 2014-05-29T06:55:00.000 600.00000

2014-05-29T11:10:00.000 -- 2014-05-29T11:15:00.000 300.00000

2014-05-29T11:30:00.000 -- 2014-05-29T11:35:00.000 300.00000

Detector issues

The detectors kept evolving this week. Brief changes are noted here, exact times and commands are on the Detector pages

D1: Rear Fast 0x60

D2: Same noise character as D3. Livetime decreases over a RHESSI day by as much as 10%. Could this be related to the light leak?

D3: Rear fast threshold change, to 0x75. Decision made based on high fast rates and correspondingly low livetime. Front 0x20.

D4: Front fast to 0xB0

D5: Front fast threshold change, to 0x60.

D6: HV decreased by 10 counts, to 1500 V.

D9: At last week's meeting we decided to keep it at ~1000V following the HV test.

shutters: Thin shutter locked in after the M event on Jun 2. D1's noise increase was blamed. Changed shutter logic to twice on D2 and twice on D5

Other notes

Shutter locked in on Jun 2

Spacecraft Management

No operations this week

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) +/- 4 minutes
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Active
Detector problems? 1, 3, 5
Personal tools